How do I graph the equation 2x - 5 = -3 + 2x?
Answer by Alan3354(69443)

How do I graph the equation 2x - 5 = -3 + 2x?
---------------
Graphs are usually done on the x-y plane, 2 dimensional.
There's no y in your equation.
----
2x - 5 = -3 + 2x?
-----
Subtract 2x from each side:
-5 = -3
Makes no sense.
Go to your next math problem, if you have one.